Given the following set of jobs waiting to run on a machine, use the Slack (least slack time remaining) dispatching rule to sequence the jobs.

Job Days Until Due Date Processing Time
A 35 30
B 27 17
C 33 30
D 41 25
A.

A,B,C,D

B.

B,C,D,A

C.

C,A,B,D

D.

B,C,A,D

E.

none of the above
Answer #1

Under the Slack (least slack time remaining) dispatching rule to sequence the jobs, we need to calculate the Slack time for each of the jobs.

Slack time remaining = Time remaining(Days) before the due date - Processing time

Job Days until the due date Processing time Slack time remaining
A 35 30 = 35 - 30 = 5
B 27 17 = 27 - 17 = 10
C 33 30 = 33 - 30 = 3
D 41 25 = 41 - 25 = 16

Hence under the Slack dispatching rule to sequence the jobs, the sequence is C, A, B, D

Hence the correct answer is option C.